OpenStudy (anonymous):

How many digits does the number \[ 2^{462} 5^{460}\] have?

OpenStudy (anonymous):

not logical

OpenStudy (anonymous):

\[\LARGE 2^{462}\cdot 5^{460}=2^{2}\cdot 2^{460}\cdot 5^{460}=4\cdot (5\cdot 2)^{460}=\] \[\LARGE =4\cdot 10^{460} \quad ??\]

OpenStudy (anonymous):

\[2^{462}5^{460}=2^{460}2^{2}5^{460}\]\[=2^{2}\left( 2*5 \right)^{460}=4\left( 10 \right)^{460}\]

OpenStudy (anonymous):

10^460*2^2

OpenStudy (anonymous):

so 461 digits?

OpenStudy (anonymous):

still 10^460 have a very big answer , her question is how many digiits would be in this question ,

OpenStudy (anonymous):

4 and "460" zeros so @dpaInc is right !

OpenStudy (anonymous):

over an infinitesimal length of digits without bypassing the speed of light ? {trollnser}

OpenStudy (anonymous):

461 is the right answer.

OpenStudy (anonymous):

@dalnc showed the analytic way, if you still need a confirmation : http://ideone.com/pwp7D